# Spatial Scale of Agglomeration and Dispersion Number, Spacing, and the Spatial Extent of Cities

### v7

@@ 2.5   General equilibrium models
where $c_{1}\equiv\frac{\mu}{\sigma-1}+\frac{\mu}{\sigma}>0$ and $c_{2}\equiv\frac{\mu^{2}}{\sigma-1}+\frac{1}{\sigma}>0$. This implies $\omega=C(\Theta)\cdot(c_{1}\Theta-c_{2}\Theta^{2})$ with $C(\Theta)\equiv 1/(1-\frac{\mu}{\sigma}\Theta-\frac{\sigma-1}{\sigma}\Theta^{2})>0$. Thus, the core trade-off between agglomeration and dispersion forces in the model is encapsulated by a concave quadratic; $c_{1}\Theta>0$ represents global agglomeration forces, and $-c_{2}\Theta^{2}<0$ captures global dispersion forces augmented by immobile demands. In this model, there is always a nonzero demand for the manufacturing goods in each region due to the presence of immobile consumers. This discourages concentration of production in one region if transport costs are high, and the uniform distribution becomes stable. As in the Braid model, agglomeration occurs if transport costs decrease from a high level.
The Tabuchi (1998) model. Tabuchi-JUE1998 extends Krugman’s model by introducing urban costs while retaining immobile consumers. In addition to the regional-scale component of Krugman, each region has an Alonso–Muth–Mills monocentric city structure. Within each region, agents commute to a central business district and face the trade-off between commuting costs and land rent. The benefit matrix reduces to
@@
where $C(\mathbf{D})$ is a matrix factor capturing general equilibrium effects, and $c_{0}=\gamma\epsilon_{1}>0$, $c_{1}=\frac{\mu}{\sigma-1}+\frac{\mu}{\sigma}>0$, and $c_{2}=\frac{\mu^{2}}{\sigma-1}\epsilon_{2}+\frac{1}{\sigma}\epsilon_{3}$. Here, $\gamma$ is the expenditure share on housing, and $\epsilon_{1},\epsilon_{2},\epsilon_{3}$ capture the effects of urban costs. For example, the larger the intraregional commuting cost parameter and/or the opportunity cost of land, the larger $\epsilon_{1}>0$ becomes, so that local dispersion force $-c_{0}$ becomes greater in magnitude. The stability of $\bar{\bm{x}}$ depends on a quadratic expression $-c_{0}+c_{1}\Theta-c_{2}\Theta^{2}$.
The agglomeration force ($c_{1}\Theta>0$) is the same as the Krugman model, while the local dispersion force ($-c_{0}<0$) is due to urban costs. The term $c_{2}\Theta^{2}$ captures global forces arising from interregional trade as well as the impacts of urban costs. The model features both local and global dispersion forces. The symmetry can be stable for both high and low levels of transport costs.
Idiosyncratic taste shocks. It is worthwhile to mention idiosyncratic preference shocks (McFadden-Book1974; McFadden-Book1978; McFadden-TRB1978) often considered in both quantitative and theoretical literature (Hunt-Simmonds-EPB1993; Waddell-JAP2002; Anas-Liu-JRS2007; Redding-Rossi-Hansberg-ARE2017; anderson1992discrete; Tabuchi-Thisse-JDE2002; Murata-JUE2003). On such taste shocks, Behrens-Murata-DP2018 demonstrated that the equilibrium condition for spatial distributions in models with idiosyncratic shocks can be equivalently represented by that in homogeneous preference models with local non-tradables markets as in Helpman. Thus, from the perspective of stabilizing forces at the symmetry, introducing idiosyncratic shocks into a model with homogeneous preference is similar to adding a negative constant to the utility gain $\omega$, i.e., to embedding an extra local dispersion force. Welfare implications are, however, different and care should be taken (Behrens-Murata-DP2018).
@@ 2.6   The three model classes
As the preceding models illustrate, the central trade-offs between agglomeration and dispersion forces often take a quadratic form in $\mathbf{D}$. That is, for all specifications of the indirect utility function $\bm{v}$ considered so far, the benefit matrix takes the form $\mathbf{V}=\mathbf{C}\cdot\big{(}c_{0}\mathbf{I}+c_{1}\mathbf{D}+c_{2}\mathbf{D}^{2}\big{)}$, with $\mathbf{C}$ and $\{c_{0},c_{1},c_{2}\}$ being model-dependent. Further, only the quadratic component determines the stability of $\bar{\bm{x}}$ in the sense that $\operatorname{sgn}\left[\omega\right]=\operatorname{sgn}\left[c_{0}+c_{1}\Theta+c_{2}\Theta^{2}\right]$ where $\operatorname{sgn}[\cdot]$ denote the sign. For brevity, we denote
$\mathbf{V}\propto c_{0}\mathbf{I}+c_{1}\mathbf{D}+c_{2}\mathbf{D}^{2},$ (22)
to indicate that such conditions are satisfied for the benefit matrix $\mathbf{V}$.
@@ C Tabuchi (Type LG)
Note: For each selected example from the three model types, the relevant quadratic component of the utility gain is plotted as a function of the transport cost index $\Theta$. The parameter values are chosen such that the stability of symmetry switches as $\Theta$ varies. In each panel, the symmetric pattern is stable for $\Theta$ such that the curve lies below the horizontal axis. In Panels (A) and (C), the intersept at $\Theta=0$ corresponds to the local dispersion force $c_{0}<0$. Local dispersion forces stabilize $\bar{\bm{x}}$ under low transport costs, and global dispersion forces under high transport costs.
@@
Model Class Stability of symmetry Spatial scale of dispersion force Examples Type L Low transport costs Local Beckmann-Book1976 Helpman-Book1998 Murata-Thisse-JUE2005 Redding-Sturm-AER2008 Allen-Arkolakis-QJE2014 Redding-Rossi-Hansberg-ARE2017, Section 3 Type G High transport costs Global Harris-Wilson-EPA1978 Krugman-JPE1991 Krugman-Venables-QJE1995 Puga-EER1999, Section 3 Forslid-Ottaviano-JoEG2003 Pfluger-RSUE2003 Type LG High and low transport costs Both local and global Tabuchi-JUE1998 Fujita-Krugman-Venables-Book1999, Section 14.4 Puga-EER1999, Section 4 Pfluger-Suedekum-JUE2008 Pfluger-Tabuchi-RSUE2010 Kucheryavyy-etal-JIE2024
Table 1: Notable examples for the three model classes
@@
Note: The section numbers in the table corresponds to those in the referenced papers. For Krugman-Venables-QJE1995 and Puga-EER1999, the share of manufacturing sector is interpreted as the spatial distribution of interest. In all the models in the table, the stability of $\bar{\bm{x}}$ hinges on the sign of a quadratic function of the form $c_{0}+c_{1}\Theta+c_{2}\Theta^{2}$ with model-dependent coefficients $\{c_{0},c_{1},c_{2}\}$. For Type L models, $c_{0}<0$ and $c_{0}+c_{1}+c_{2}>0$, so that $\bar{\bm{x}}$ is unstable for large $\Theta$. Some Type L models incorporate market crowding effects across regions, meaning that $c_{2}$ can be negative. However, for any values of structural parameters that allow for multiple equilibria, the symmetry is unstable when $\Theta$ is large. For Type G models, $c_{0}\geq 0$, indicating that there is no net local dispersion forces.
@@
Under the quadratic representation in Eq. 22, we can define the three prototypical model classes mentioned in Section 1. Specifically, we classify models depending on the conditions under which symmetry can be stable: high transport costs (large $\Theta$), low transport costs (small $\Theta$), or both. We refer to the three model classes as Type L, Type G, and Type LG, where L denotes “local” and G denotes “global.” We can summarize as follows, with $\{c_{k}\}$ being model-dependent constants: [^13] See Section A.2 for the formal definitions. To classify models, we assume that agglomeration forces are sufficiently strong so that the stability of the symmetric equilibrium depends on the transport cost index $\Theta$, meaning that $\omega$ cuts the horizontal axis at some $\Theta$. This excludes the case in which the symmetry is stable (or unstable) for all $\phi$.
@@
  • Type L emphasizes local dispersion forces. The symmetry is stable if transport costs are low, and unstable if transport costs are high. A typical form is $\mathbf{V}\propto c_{0}\mathbf{I}+c_{1}\mathbf{D}$ with $c_{0}<0$ indicating the local dispersion force.
@@
  • Type G emphasizes global dispersion forces. The symmetry can be stable if transport costs are high, and unstable if transport costs are low. Typically, $\mathbf{V}\propto c_{1}\mathbf{D}+c_{2}\mathbf{D}^{2}$, with $c_{2}<0$ representing the global dispersion force.
@@
  • Type LG is the synthesis of Type G and Type L, and has both local and global dispersion forces: $\mathbf{V}\propto c_{0}\mathbf{I}+c_{1}\mathbf{D}+c_{2}\mathbf{D}^{2}$, where $c_{0},c_{2}<0$ and $c_{1}>0$. The symmetry can become stable for high and low transport cost levels.
Figure 4 illustrates typical forms of the relevant quadratic component of the utility gain for the three model classes. Table 1 lists notable examples in the literature.
In the quadratic representation in Eq. 22, the coefficients $\{c_{0},c_{1},c_{2}\}$ are functions of the model’s structural parameters, excluding transport costs encapsulated in $\mathbf{D}$. The constant $c_{0}$ captures all the forces within each location, i.e., the local forces. The latter two coefficients $c_{1}$ and $c_{2}$ represent the global forces across locations, with different orders of $\mathbf{D}$. The first-order coefficient $c_{1}$ reflects the direct effects from one region to another, such as the agglomeration externalities in the Beckmann model. Similarly, $c_{2}$ captures indirect spillovers, such as the market crowding effect at intermediate regions, as seen in gravity-type structure of interregional trade.
Each coefficient summarizes the composite effect at the corresponding order of $\mathbf{D}$, and its sign indicates the net impact on $\omega$. For instance, suppose that there are both local agglomeration economies and diseconomies. Then, $c_{0}>0$ indicates that local agglomeration economies dominate diseconomies, and vice versa.
@@ 3   Many regions
Our analysis so far has highlighted that different dispersion forces become consequential at distinct transport cost regimes. Although this insight has been recognized in various forms in the theoretical literature, its implications on spatial patterns for settings with more than two regions remain underexplored. This section considers an $N$-region economy. All variables and functions such as $\bm{x}$ or $\bm{v}$ are straightforwardly extended.
We must define the “fully dispersed” state to study spatial agglomeration. A natural candidate in an $N$-region setting is the uniform spatial distribution $\bar{\bm{x}}\equiv(\frac{1}{N},\frac{1}{N},\ldots,\frac{1}{N})$. To employ this distribution as the analytical ground, we assume a symmetric geography following krugman1993number, in which the regions are homogeneous and equidistantly placed over a circle, and transport is possible only along the circumference:
Figure 5: $N$-region symmetric circle.
@@
The proximity matrix is given by $\phi_{ij}=\phi^{\ell_{ij}}$, where $\phi\in(0,1)$ is the ease of transport between two consecutive regions, and $\ell_{ij}\equiv\min\{|i-j|,N-|i-j|\}$ is the distance between regions $i$ and $j$ over the circumference. All regions are symmetric regarding their local fundamentals (e.g., innate amenity or productivity). ∎
Figure 5 illustrates the geography. Since every region has the same level of accessibility to the other regions in a circle, Assumption C abstracts away geographical advantages (Matsuyama-RIE2017), the advantages from each region’s unique position within transport networks. Combined with the perfect symmetry in other regional fundamentals, [^14] See Assumption S in Section A.2 for the formal version of this part of Assumption C. the uniform distribution $\bar{\bm{x}}$ is always a spatial equilibrium.
3.1   Benefit matrix and the stability of symmetry
If the symmetric equilibrium does not withstand small migration shocks, it is unstable and some form of agglomeration should occur. A migration shock can be represented as $\bar{\bm{x}}+\bm{z}$, where $\bm{z}=(z_{i})_{i\in\mathcal{I}}$ is a deviation pattern such that $\sum_{i\in\mathcal{I}}z_{i}=0$. For each region $i$, $z_{i}>0$ represents a population increase, and $z_{i}<0$ a decrease. Clearly, the $N$-region setting allows for a wider range of deviation patterns than the two-region setting. [^15] Specifically, the only deviation pattern allowed in the two-region case is $\bm{z}=(\epsilon,-\epsilon)$ with small $\epsilon$.
To assess the local stability of the full dispersion $\bar{\bm{x}}$, we examine whether any infinitesimal deviation pattern can improve the utility of movers. The following lemma shows that the benefit matrix $\mathbf{V}$ once again plays a central role:

### v8

@@ 2.4  General equilibrium models
where $c_{1}\equiv\frac{\mu}{\sigma-1}+\frac{\mu}{\sigma}>0$ and $c_{2}\equiv\frac{\mu^{2}}{\sigma-1}+\frac{1}{\sigma}>0$. The gain function is then $\omega=C(\Theta)\cdot(c_{1}\Theta-c_{2}\Theta^{2})$ with $C(\Theta)\equiv(1-\frac{\mu}{\sigma}\Theta-\frac{\sigma-1}{\sigma}\Theta^{2})^{-1}>0$, where the core trade-off is captured by $\omega^{\sharp}\equiv c_{1}\Theta-c_{2}\Theta^{2}$. The first term $c_{1}\Theta>0$ represents global agglomeration forces, and the second term $c_{2}\Theta^{2}<0$ captures global dispersion forces augmented by immobile demands. As in the Braid model, dispersion is preferred if transport costs are high (small $\phi$), and agglomeration occurs otherwise (large $\phi$).
The Tabuchi-JUE1998 model. Tabuchi integrated urban costs into Krugman’s framework with immobile consumers. In addition to the regional-scale component of Krugman, each region has an Alonso–Muth–Mills monocentric city structure. Within each region, agents commute to a central business district and face the trade-off between commuting costs and land rent. The benefit matrix for the Tabuchi model reduces to
@@
where $C(\mathbf{D})$ is a matrix factor that captures general equilibrium effects, and $c_{0}=\gamma\epsilon_{1}>0$, $c_{1}=\frac{\mu}{\sigma-1}+\frac{\mu}{\sigma}>0$, and $c_{2}=\frac{\mu^{2}}{\sigma-1}\epsilon_{2}+\frac{1}{\sigma}\epsilon_{3}$. Here, $\gamma$ is the share of housing expenditures, and $\epsilon_{1},\epsilon_{2},\epsilon_{3}$ capture the effects of urban costs. For example, the larger the commuting cost parameter and/or the opportunity cost of the land, the larger $\epsilon_{1}>0$ becomes, so that the local dispersion force $c_{0}$ becomes more pronounced. The stability of $\bar{\bm{x}}$ depends on the quadratic expression $\omega^{\sharp}\equiv-c_{0}+c_{1}\Theta-c_{2}\Theta^{2}$.
The agglomeration force ($c_{1}\Theta>0$) is the same as the Krugman model, while $c_{0}<0$ captures the local dispersion force due to urban costs. The term $c_{2}\Theta^{2}$ captures net global forces that include the impacts of urban costs through general equilibrium. The model features both local and global dispersion forces, and the symmetric equilibrium can be stable for both high and low levels of transport costs.
Idiosyncratic taste shocks. Idiosyncratic preference shocks (McFadden-Book1974; McFadden-Book1978; McFadden-TRB1978) are important both in the quantitative and theoretical literature (Hunt-Simmonds-EPB1993; Waddell-JAP2002; Anas-Liu-JRS2007; Redding-Rossi-Hansberg-ARE2017; anderson1992discrete; Tabuchi-Thisse-JDE2002; Murata-JUE2003). Regarding taste shocks, Behrens-Murata-DP2018 demonstrated that the spatial equilibrium condition in models with idiosyncratic shocks can be equivalently represented by that in homogeneous preference models with local non-tradables markets as in Helpman-Book1998. Thus, from the perspective of stabilizing forces at $\bar{\bm{x}}$, introducing idiosyncratic shocks into a model with homogeneous preference is equivalent to adding a negative constant to the utility gain $\omega$, i.e., to embedding an additional local dispersion force. Welfare implications are, however, different and care should be taken (Behrens-Murata-DP2018).
@@ 2.5  The three model classes
As the above examples illustrate, for a wide class of spatial models,
$\mathbf{V}=C(\mathbf{D})\cdot\bigl(c_{0}\mathbf{I}+c_{1}\mathbf{D}+c_{2}\mathbf{D}^{2}\bigr),\quad\text{and hence}\quad$ (29)
$\omega=C(\Theta)\cdot\bigl(c_{0}+c_{1}\Theta+c_{2}\Theta^{2}\bigr),$ (30)
where $C$ and the coefficients $\{c_{0},c_{1},c_{2}\}$ are model dependent. In this representation, the stability of $\bar{\bm{x}}$ is governed by the quadratic term, since the sign of $\omega$ is entirely determined by the sign of $c_{0}+c_{1}\Theta+c_{2}\Theta^{2}$.
@@ C Tabuchi (Type LG)
Note: For selected examples from the three model types, the “net” utility gain is plotted as a function of the proximity gain $\Theta$. Larger $\Theta$ corresponds to smaller $\phi$. In each panel, $\bar{\bm{x}}$ is stable for $\Theta$ such that the curve lies below the horizontal axis. The parameter values are chosen such that the stability of $\bar{\bm{x}}$ depends on $\Theta$. Local dispersion forces stabilize $\bar{\bm{x}}$ for small $\Theta$, and global dispersion forces for large $\Theta$.
@@
Model Class Stability of $\bar{\bm{x}}$ Dominant dispersion force Examples Type L Low transport costs Local Beckmann-Book1976 Helpman-Book1998 Murata-Thisse-JUE2005 Redding-Sturm-AER2008 Allen-Arkolakis-QJE2014 Redding-Rossi-Hansberg-ARE2017, Section 3 Type G High transport costs Global Harris-Wilson-EPA1978 Krugman-JPE1991 Krugman-Venables-QJE1995 Puga-EER1999, Section 3 Forslid-Ottaviano-JoEG2003 Pfluger-RSUE2003 Type LG High and low transport costs Both local and global Tabuchi-JUE1998 Fujita-Krugman-Venables-Book1999, Section 14.4 Puga-EER1999, Section 4 Pfluger-Suedekum-JUE2008 Pfluger-Tabuchi-RSUE2010 Kucheryavyy-etal-JIE2024
Table 1: Notable examples
@@
Note: The section numbers in the table correspond to those in the referenced papers. For Krugman-Venables-QJE1995 and Puga-EER1999, the spatial distribution of interest is the share of manufacturing sector. In all the models in the table, the stability of $\bar{\bm{x}}$ hinges on the sign of a quadratic function of the form $c_{0}+c_{1}\Theta+c_{2}\Theta^{2}$ with model-dependent coefficients $\{c_{0},c_{1},c_{2}\}$.
@@
For models of this form, we can define three prototypical classes according to the transport cost conditions under which the symmetric equilibrium is stable: low transport costs (small $\Theta$ or large $\phi$), high transport costs (large $\Theta$ or small $\phi$), or both. For convenience, we refer to the three model classes as Type L, Type G, and Type LG, where L and G stand for “local” and “global,” respectively. [^12] See Appendix A.3 for formal definitions. Here, we restrict our attention to cases in which the agglomeration forces are neither too weak nor too strong, so that the symmetric equilibrium is stable for some values of $\phi$ and unstable for others.
With $\{c_{k}\}$ being model-dependent constants, we can summarize as follows:
@@
  • Type L emphasizes local dispersion forces as the dominant dispersion mechanism. The symmetric equilibrium is stable if transport costs are low, and unstable if transport costs are high. $\mathbf{V}=C(\mathbf{D})\cdot(c_{0}\mathbf{I}+c_{1}\mathbf{D})$ with $c_{0}<0$ indicating the local dispersion force.
@@
  • Type G emphasizes global dispersion forces as the dominant dispersion mechanism. The symmetric equilibrium is stable if transport costs are high, and unstable if transport costs are low. $\mathbf{V}=C(\mathbf{D})\cdot(c_{1}\mathbf{D}+c_{2}\mathbf{D}^{2})$, with $c_{2}<0$ representing the global dispersion force.
@@
  • Type LG has both local and global dispersion forces as the dominant dispersion mechanisms. $\mathbf{V}=C(\mathbf{D})\cdot(c_{0}\mathbf{I}+c_{1}\mathbf{D}+c_{2}\mathbf{D}^{2})$, where $c_{0},c_{2}<0$ and $c_{1}>0$. The symmetric equilibrium is stable for both high and low transport cost levels.
Figure˜5 illustrates the typical shapes of the quadratic term $c_{0}+c_{1}\Theta+c_{2}\Theta^{2}$ in each class, and Table˜1 lists representative examples.
The model-dependent coefficients $c_{0},c_{1},c_{2}$} are functions of the model’s structural parameters, with transport costs entering only through $\mathbf{D}$. The constant $c_{0}$ captures the forces operating within regions and therefore represents the local component. The coefficients $c_{1}$ and $c_{2}$ represent the forces that operate across regions. The first-order term $c_{1}$ reflects direct interregional effects, such as agglomeration spillovers in the Beckmann model. The quadratic term $c_{2}$ captures higher-order spillovers. [^13] In many-region settings, $c_{2}$ reflects indirect interactions mediated by third regions. For example, in models with interregional trade, agents in the region $i$ are affected by the population of the region $j$ because agents in the region $i$ compete with those in the region $j$ for income generated in other regions $k=1,2,\ldots$. Such effects typically arise second order in transport frictions, as they depend on the transport costs between regions $i$ and $k$, as well as between regions $j$ and $k$.
Each coefficient $c_{k}$ in the $c_{0},c_{1},c_{2}$} representation captures the composite general equilibrium effect associated with the corresponding order of $\mathbf{D}$. Its sign therefore reflects the net contribution to $\omega$ in the order of $\Theta$. For example, if there are both local agglomeration economies and diseconomies, the sign of $c_{0}$ reveals which force dominates: $c_{0}>0$ indicates net agglomeration, while $c_{0}<0$ indicates net dispersion. In the Allen–Arkolakis framework, for example, one obtains $c_{0}=\alpha-\beta-\frac{1+\alpha}{\sigma}$, and $\alpha<\beta$ is a sufficient condition for the net local effect $c_{0}$ to be negative.
@@ 3  Many regions
This section examines how the proposed taxonomy of spatial models maps to the endogenous spatial patterns and their comparative statics in an $N$-region economy. All variables and functions (e.g., $\bm{x}$, $\bm{v}$, $\phi_{ij}]$, $\mathbf{D}$) are straightforwardly extended. The set of regions is now denoted by $\mathcal{I}\equiv\{1,2,\ldots,N\}$.
We focus on a stylized geography in which homogeneous regions are symmetrically placed over a circle and transport is possible only along the circumference (Fig.˜6).
Figure 6: NN-region symmetric circle.
@@
The proximity matrix is given by $\phi_{ij}=\phi^{\ell_{ij}}$, where $\phi\in(0,1)$ is the ease of transport between two consecutive regions, and $\ell_{ij}\equiv\min\{|i-j|,N-|i-j|\}$ is the distance between regions $i$ and $j$ over the circumference. All regions are symmetric regarding their local fundamentals (e.g., innate amenity or productivity). In addition, $N$ is a multiple of four. [^14] This restriction on $N$ is only for expositional simplicity. See Remark 3 in Appendix A.
This abstracts away the advantages from each region’s unique geographic position: every region has the same level of geographic accessibility in a circle. Combined with the perfect symmetry in other regional fundamentals, the symmetric distribution $\bar{\bm{x}}\equiv(\frac{1}{N},\frac{1}{N},\ldots,\frac{1}{N})$ is always a spatial equilibrium (Lemma˜3 in Appendix A.3).
3.1  The stability of the symmetric equilibrium
Endogenous agglomeration occurs when the symmetric equilibrium $\bar{\bm{x}}$ is unstable, that is, when it does not withstand small migration shocks. In an economy with $N$ regions, such shocks can be represented as
$\bar{\bm{x}}+\epsilon\bm{z}=(\bar{x}+\epsilon z_{1},\bar{x}+\epsilon z_{2},\ldots,\bar{x}+\epsilon z_{N})$ (31)
where $\epsilon$ is a sufficiently small scalar. The vector $\bm{z}=(z_{i})_{i\in\mathcal{I}}$ is a deviation pattern: $z_{i}>0$ indicates an inflow into region $i$ and $z_{i}<0$ an outflow. Since the total population is fixed, we require $\sum_{i\in\mathcal{I}}z_{i}=0$. The $N$-region setting admits a substantially richer set of deviation patterns than the two-region case, because $(\bar{x}+\epsilon,\bar{x}-\epsilon)$ is the only possible perturbation in the two-region setting. Below, we normalize $\|\bm{z}\|=1$.
Analogous to the two-region case, the expected utility gains for marginal movers under migration shocks is closely related to the benefit matrix $\mathbf{V}\equiv\big[\frac{\partial{v_{i}}}{\partial{x_{j}}}(\bar{\bm{x}})\big]_{i\in\mathcal{I},j\in\mathcal{I}}$.